# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract
Recent advancements in moiré engineering motivate study of the behavior of strongly-correlated electrons subject to substantial orbital magnetic flux. We investigate the triangular lattice Hofstadter-Hubbard model at one-quarter flux quantum per plaquette and a density of one electron per site, where geometric frustration has been argued to stabilize a chiral spin liquid phase intermediate between the weak-coupling integer quantum Hall and strong-coupling 120 antiferromagnetic phases. In this work, we use Density Matrix Renormalization Group methods and analytical arguments to analyze the compactification of the Hofstadter-Hubbard model to cylinders of finite radius. We introduce a glide particle-hole symmetry operation which for odd-circumference cylinders, we show, is spontaneously broken at the quantum Hall to spin liquid transition. We further demonstrate that the transition is associated with a diverging correlation length of a charge-neutral operator. For even-circumference cylinders the transition is associated with a dramatic quantitative enhancement in the correlation length upon threading external magnetic flux. Altogether, we argue that the 2+1D CSL-IQH transition is in fact continuous and features critical correlations of the charge density and other spin rotationally-invariant observables.
@@
Introduction—The Kalmeyer-Laughlin chiral spin liquid (CSL) is one of the earliest and most important examples of topological order [1, 2, 3] and has been extensively studied in spin systems, both numerically [4] and through partonic mean-field approaches [5, 6, 7]. Recently, considerable efforts have been devoted to identifying electronic settings in which the CSL may emerge to connect more directly to solid state experiments, in which electrons—rather than spins—are the fundamental degrees of freedom and may be far from their Mott limit. These systems enable the exploration of the Mott transition of strongly-interacting electrons in strong magnetic fields, a regime that is becoming experimentally accessible both at non-zero [8, 9] and zero Zeeman coupling [10, 11]. They also set the stage for the study of charge fluctuation-driven transitions out of the CSL phase into other phases such as superconductors [12], exotic charge-density waves [13], or quantum Hall states [14, 11].
@@
Figure 1: (a) Upper panel: The Hofstadter-Hubbard model on the 2D triangular lattice, with magnetic translations $T_{x,y}$, on-site Hubbard $U$ interaction, and flux per triangular plaquette $\Phi_{\triangle}$ indicated. Lower left panel: “Hofstadter butterfly” representation of spectrum of non-interacting model in the plane of single-particle energy $E$ and flux per plaquette $\Phi_{\triangle}$. $\Phi_{\triangle}=\pi/2$ is the unique flux for which the Hamiltonian enjoys a doubled “magnetic” unit cell and a particle-hole symmetry $\mathcal{P}$ relating opposite $C=\pm 1$ bands [lower right panel]. (b) A finite segment of the infinite-length, circumference-$L_{y}$ cylinder geometry, threaded by $\Phi_{\text{ext}}$ external magnetic flux. With the chosen “YC” cylindrical geometry, the translation $T_{y}$ is orthogonal to the cylinder axis and $x\in\mathbb{Z}$ indexes rings. (c) Expected 2+1D phase diagram at half-filling. As interactions increase, the spin-degenerate integer quantum Hall (IQH) ground state undergoes a quantum phase transition into a Kalmeyer-Laughlin chiral spin liquid (CSL). By studying both even and odd finite-circumference cylinders, we shed light on the existence and nature of this transition.
@@
A model electronic system thought to host both conventional and CSL electronic phases is the triangular Hofstadter-Hubbard model (see Fig. 1a) in the regime of strong orbital magnetic flux, where the CSL may be favored energetically by both geometric frustration [15, 16] and the explicit breaking of time-reversal symmetry [17, 7]. The Hofstadter-Hubbard model is parametrized by an on-site interaction $U$ and a nearest neighbor hopping with Peierls phases corresponding to a flux of $\Phi_{\triangle}$ per elementary plaquette. When $\Phi_{\triangle}=\pi/2$ per triangle, this system realizes a spin-unpolarized integer quantum Hall (IQH) insulator at small values of the Hubbard interaction, with $\sigma_{xy}=2e^{2}/h$, obtained by filling the lower topological sub-band in both spin sectors. This state possesses a gap to spin and charge excitations, is topological and has a spin-degenerate charged edge mode. At very large $U$ and density of one electron per site, the state is a Mott insulator: the leading term in the $t/U$ expansion around the infinite-$U$ Mott limit shows that the low-energy physics is governed by the conventional antiferromagnetic nearest-neighbor $S=1/2$ Heisenberg model with $J_{H}\sim t^{2}/U$, which has $120^{\circ}$ antiferromagnetic order at $T=0$, so that the ground state is non-topological, with fully-gapped charge excitations and soft (Goldstone) spin excitations.
@@
Carrying out the $t/U$ expansion to higher order suggests the existence of a CSL phase intermediate between the Mott/AF and IQH phases: the next correction is a third-order chiral term $S\cdot S\times S$ with coefficient $J_{\triangle}\sim t^{3}\sin\Phi_{\triangle}/U^{2}$. In the presence of the magnetic flux this term explicitly breaks both time-reversal and parity symmetry [17, 7]. Numerical studies of the $J_{H}$-$J_{\triangle}$ spin Hamiltonian have indeed revealed a CSL ground state for large enough $J_{\triangle}/J_{H}$ [18, 19], although at these values the accuracy of truncating the $t/U$ expansion at third order may be questioned. Very recently Kuhlenkamp et al. presented cylinder iDMRG evidence of an intermediate CSL phase for $U\gtrsim 10t$ at fluxes $\Phi_{\triangle}=\pi/3,\pi/2$ [11]. The CSL phase was identified by the absence of charge pumping as flux was threaded through the cylinder and the presence of appropriate degeneracies in the entanglement spectrum.
Supposing one accepts the existence of an intermediate CSL phase, the question of the nature of the IQH-CSL transition arises. A parton theory in which one decomposes the electron into a bosonic chargon $b$ carrying the charge and a fermionic spinon $f$ carrying the spin index, $c_{\sigma}=bf_{\sigma}$ has been proposed [14, 12]. This decomposition requires an emergent gauge field $a$, corresponding to the relative phase between the partons, under which they carry opposite unit $U(1)$ charge. In both the IQH and CSL phases the spinon occupies both $C=1$ magnetic sub-bands descending from the electronic bandstructure, while at the electronic IQH-CSL phase transition the chargon and the gauge field together undergo a superconductor-Mott phase transition [12]. The dynamical gauge field $a$ is responsible not only for producing the correct topological response on both sides of the transition, but also for gapping out the charge sector in the IQH phase through the Anderson-Higgs mechanism [20, 21], and allowing for spinon deconfinement in the CSL [22, 23, 24]. The parton representation and related slave-particle numerical studies on honeycomb and square lattices  [14, 25] have provided a physically intuitive, yet incomplete mean-field picture of the putative continuous transition. The iDMRG investigation of Ref. [11] revealed the IQH-CSL transition only as a crossover indicated by a weak maximum in a correlation length. Further investigation of the CSL phase and the IQH-CSL transition is therefore needed.
In this work we use a compactification to an infinite-length, finite-radius cylinder [see Fig. 1(b)] in combination with a symmetry analysis with large-scale iDMRG simulations [26, 27] on cylinders of circumference $L_{y}=3$ to $6$ to characterize the nature of the IQH and CSL states, as well as the transition between them. Our symmetry analysis is based on an operator that is a combination of particle-hole interchange and translation, which we show explicitly is spontaneously broken for interactions above a critical interaction strength for odd-circumference cylinders. We further show that the broken symmetry occurs together with the degeneracy pattern in the entanglement spectrum expected for a CSL phase. We present numerical evidence that this transition is in fact continuous and argue that in the limit of infinite cylinder radius, this transition evolves into a continuous 2+1D IQH-CSL quantum phase transition featuring critical fluctuations in the charge density and other spin rotationally-invariant observables.
@@
Figure 2: CSL symmetry-breaking at odd circumference. (a) Graphical illustration of the action of the particle-hole symmetry operator $\mathcal{P}$ which transforms the gauge flux through a ring of an odd-$L_{y}$ cylinder, from $\phi$ to $\pi-\phi$. (b) Upper panel: In an odd-$L_{y}$ CSL ground state, the alternating pattern $+\pi/2,-\pi/2,\dots$ of these gauge fluxes spontaneously breaks $\mathcal{P}T_{x}$ symmetry, whose action is equivalent to the threading of $\pi$ gauge flux. Lower panel: the low-lying entanglement eigenvalues at cuts between two pairs of rings, which not only exhibit the approximate degeneracy expected for the CSL at each $S_{z}$ [see circled numerals], but also transform as half-integer spin representations only at every other cut, signalling SSB of $\mathcal{P}T_{x}$. (c) In contrast, the internal gauge flux configuration [upper panel] and entanglement spectrum [lower panel] of the IQH state are consistent with it being a unique, symmetry-preserving ground state.
@@
Model, symmetry operator and cylinder compactification—We study the Hofstadter-Hubbard model on the triangular lattice, defined pictorially in Fig. 1(a). In this model, electrons hop between nearest-neighboring sites and experience an Aharonov-Bohm phase of $\Phi_{\triangle}$ upon orbiting a triangular plaquette. Note that this model focuses on the consequences of the orbital motion, and sets the Zeeman coupling to zero. We fix $\Phi_{\triangle}=\pi/2$, which endows the Hamiltonian with a unitary particle-hole (PH) symmetry $\mathcal{P}$ acting as [28]
@@
on fermionic operators, where we have attached a conventional spin rotation $\sigma_{y}$ to the naive definition of particle-hole transformation in order for $\mathcal{P}$ to leave the spin operator $\bm{S}=\frac{1}{2}c^{\dagger}\bm{\sigma}c$ invariant (see SM for details [29]). The flux induces a magnetic translation algebra $T_{x}T_{y}=-T_{y}T_{x}$ that we represent in a gauge where
$T_{x}c^{\dagger}_{x,y}T_{x}^{-1}=(-1)^{y}c^{\dagger}_{x+1,y},\quad T_{y}c^{\dagger}_{x,y}T_{y}^{-1}=c^{\dagger}_{x,y+1}.$ (2)
@@
The commuting translations $(T_{x})^{2},T_{y}$ define a larger “magnetic” unit cell, whose resulting two sub-bands are related by $\mathcal{P}$ and carry opposite Chern number $C=\pm 1$. We fix the filling to $n=1$ electron per lattice site, which in the absence of interactions corresponds to full filling of the lowest spin-degenerate magnetic sub-band.
We now present the dimensional reduction of the problem to a cylinder. The key results of our study rely on the different symmetry breaking patterns that occur for even and odd circumferences $L_{y}$, which stem from generic properties of magnetic fields (both external and internal) in the chosen geometry. To illustrate how this arises, let us first consider the closed “tin can”-shaped surface specified by a pair of adjacent rings at $x$ and $x+1$. This surface is the union of $2L_{y}$ triangles on the outside of the cylinder, as well as an $L_{y}$-sided polygonal surface at $x$ and another at $x+1$ whose edges are the rings. Since $\pi/2$ external flux penetrates each outer triangle, Gauss’s law requires that the external magnetic flux through the two adjacent rings differ by $\pi L_{y}$ [30, 31]. These ring fluxes are therefore uniform when $L_{y}$ is even, but are staggered when $L_{y}$ is odd, which explicitly breaks $T_{x}$ symmetry.
Let us momentarily focus on the odd case and fix the sequence of ring fluxes to alternate between 00 and $\pi$. In this case, the system nonetheless respects a combined translation-like symmetry, namely the product of $\mathcal{P}$ and $T_{x}$—which we refer to as PH-glide—in spite of each being individually broken by the compactification. The IQH state, being adiabatically connected to the (unique) ground state of the non-interacting Hamiltonian, is invariant under PH-glide symmetry.
We argue, however, that the CSL phase must break $\mathcal{P}T_{x}$ spontaneously when $L_{y}$ is odd. To see this, note that the CSL is a Mott insulating phase. Thus, we may consider the limit where charge fluctuations are frozen out, e.g. by increasing $U\to\infty$ while holding the three-spin chiral term at a fixed favorable strength. Since the particle-hole operation is transparent to the spin operator, $\mathcal{P}\bm{S}\mathcal{P}^{-1}=\bm{S}$, this results in a spin model with a pure translation symmetry $t_{x}$.00footnotetext: We note a similar observation in a simpler 1+1D model [32]. Since each ring contains an odd number of spin-$1/2$ moments, the Lieb-Schulz-Mattis theorem requires that any gapped ground state be at least two-fold degenerate [33], e.g., by spontaneously breaking $t_{x}$. One way to understand the translation symmetry breaking of the CSL, in particular, is by leveraging its anyonic content. Namely, the properties of the CSL under translations are captured by viewing it has hosting a spinon in each unit cell of the triangular lattice [34, 35, 36]. Because these spinons have semionic statistics, the action of $t_{x}$ translation—which moves $L_{y}$ many spinons across any given cut—necessarily permutes the two minimally-entangled CSL ground states [37] when $L_{y}$ is odd [38, 39]. In other words, the only $t_{x}$-symmetric ground states are long-range-entangled CSL “cat” states, implying that translation is spontaneously broken [40].
Figure 3: Evidence for the IQH-CSL phase transition. (a) Upper panel: Sketch of three rings of cylinder at indicated $x$ values, with solid and dashed lines showing along which staggered-Heisenberg order parameter $\mathcal{O}_{\delta\mathrm{Heis}}$ [defined in Eq. (3)] changes under glide particle-hole symmetry $\mathcal{P}T_{x}$ on odd-circumference cylinders. Lower panel: Magnitude of the staggered Heisenberg order parameter plotted against $U/t$ for cylinder widths indicated and with $\mathcal{P}T_{x}$ preserving and broken phases indicated by the check and cross symbols. Consistent with the recovery of this symmetry at the putative 2+1D critical point, the magnitude of the order parameter [bottom left] decreases substantially from $L_{y}=3$ to $5$. (b-d) Correlation length in the $(Q,S_{z},k_{y})=\bm{0}$ charge sector, plotted against interaction strength, for a range of MPS bond dimensions [increasing light to dark] and extrapolated to infinite bond dimension [gray points]; the insets specify the external flux threading each of the cylinder rings. (b) At circumference $L_{y}=5$, the $\mathcal{P}T_{x}$ symmetry-breaking is associated with a diverging correlation length. (c) While both the IQH and CSL states are fully-symmetric on even-$L_{y}$ cylinders, the dominant correlation length is consistently charge-neutral, peaking strongly for $\Phi_{\text{ext}}=\pi/2$ external flux threading the cylinder (see inset). The maximum extrapolated value of the correlation length increases from $L_{y}=4$ to $6$, consistent with an approach to a continuous transition in the thermodynamic limit.
It is illustrative to phrase the above discussion in terms of operations on the internal gauge field introduced in the parton construction. For simplicity, let us focus on saddle point configurations of the gauge field on each side of the transition, ignoring fluctuations. In both the CSL and IQH ground states, the gauge flux through the triangular surface plaquettes should match the external flux, $\phi_{\triangle}=\pi/2$ [6, 7, 12]. Thus, the only remaining gauge-invariant degree of freedom is the gauge flux through the cylinder rings; assuming $T_{x}^{2}$ is unbroken and $L_{y}$ is odd, we may parameterize the two fluxes as $(\phi,\phi+\pi)$, differing by $\pi$ due to Gauss’s law, as invoked above. Since the minimally-entangled CSL ground states are adiabatically connected to the Mott limit in which $\mathcal{P}$ acts trivially, their ring flux patterns are restricted to be $(\pi/2,-\pi/2)$ and $(-\pi/2,\pi/2)$, which are insensitive to the action of $\mathcal{P}$ on ring fluxes, namely $\phi\to\pi-\phi$ (Fig. 2a). We readily see that these two configurations are related by $\mathcal{P}T_{x}$, exposing the two-fold degeneracy and SSB of the CSL (Fig. 2b, top) [41]. In the IQH phase, on the other hand, the Meissner effect pins together the internal and external gauge fields [42]. The resulting flux pattern $(0,\pi)$ maps into itself under $\mathcal{P}T_{x}$, consistent with a unique, gapped ground state (Fig. 2c, top).
While the small $U$ IQH phase does not break the glide-PH symmetry and the intermediate $U$ CSL phase has a two-fold symmetry breaking pattern on odd circumference cylinders, the large $U$ $120^{\circ}$ phase breaks translational symmetry (including the glide PH) in a threefold spatial pattern.
Numerical observation of phase transition—We first comment on evidence for the phases themselves. The primary diagnostic of chiral topological behavior is the behavior of the charge-resolved entanglement spectrum (ES) [43, 44] at a bipartition of the system into two semi-infinite halves, in this case between any pair of adjacent rings [40, 11]. We exploit the relationship between the physical and entanglement edges [45, 43, 46], namely that the low-lying states in the entanglement spectrum correspond to those of the edge CFT [47, 48, 49, 50].
Within the bulk-boundary correspondence, the CSL is expected to host gapless edges described at low energies by a free bosonic conformal field theory [51]. The counting of low-energy states as a function of momentum in each conserved charge sector can be obtained combinatorially: for fixed $n\in\mathbb{Z}_{\geq 0}$ and associated momentum $q_{y}=2\pi n/L_{y}$, one considers a linearly dispersing mode and imagines exciting states from below to above the Fermi level. Note that these bosonic objects are charge and spin-neutral but carry momentum. The counting sequence is namely the number of integer partitions of $n$: $(1,1,2,3,5,\dots).$ On the other hand, the IQH phase has two independent bosonic edge modes. Here, the total momentum must be the sum of those of the two flavors. Working out the “two-color” integer partitions of $n$ yields the counting $(1,2,5,10,20\dots).$ These expectations are borne out in our numerics; see Fig. 2(b,c).
We now investigate the transition. Consistent with the expectations above, we confirm that the CSL phase spontaneously breaks $\mathcal{P}T_{x}$ symmetry on odd-$L_{y}$ cylinders. We first find that the bipartite ES in the CSL phase exhibits a two-ring periodicity, which implies SSB. In contrast, the ES of the IQH is fully invariant under the action of $\mathcal{P}T_{x}$, as shown in Fig. 2(c). Second, by scanning $U/t$ and inspecting the transfer matrix spectrum, we identify a diverging correlation length for the accessible odd-circumference cylinder geometries, namely $L_{y}=3,5$. The divergence is found to belong to the $(Q,S_{z},k_{y})=\bm{0}$ charge sector (see Fig. 3b for the $L_{y}=5$ and the SM [29] for $L_{y}=3$). The spontaneous breaking of glide-PH symmetry is the source of this quasi-(1+1)D criticality. To substantiate this, we identified an order parameter $\mathcal{O}_{\delta\mathrm{Heis}}=\sum_{m}\mathcal{O}^{\delta\mathrm{Heis}}_{2m}$, where
@@
The operator $\mathcal{O}_{\delta\mathrm{Heis}}$ is both charged under $\mathcal{P}T_{x}$ and condenses in the CSL phase [see Fig. 3(a)], signaling SSB. We remark that threading external flux $\Phi_{\text{ext}}=\pi/2$ through an odd-circumference cylinder reduces the transition to a crossover, owing to the fact that flux threading explicitly breaks glide-PH symmetry unless $\Phi_{\text{ext}}\in\pi\mathbb{Z}$ (see SM [29]).

### v2

@@ Abstract
Recent advances in moiré engineering motivate the study of lattice models of strongly-correlated electrons subjected to substantial orbital magnetic flux. We analyze the triangular lattice Hofstadter-Hubbard model at one-quarter flux quantum per plaquette and a density of one electron per site, where a chiral spin liquid phase may exist between weak-coupling integer quantum Hall and strong-coupling 120 antiferromagnetic phases. We use matrix product state methods and analytical arguments to investigate this model compactified to cylinders of finite circumference. We uncover a glide particle-hole symmetry operation which, we argue, is spontaneously broken at the quantum Hall to spin liquid transition on odd-circumference cylinders. We numerically verify the spontaneous symmetry breaking and further demonstrate that this transition is associated with algebraic long-range correlations of various spin-singlet, charge-neutral operators. For even-circumference cylinders, the transition becomes a crossover associated with a large correlation length that grows substantially with circumference. Our findings suggest that in the two-dimensional limit, the transition to a chiral spin liquid phase is continuous and features critical fluctuations of the current.
@@
Introduction—The Kalmeyer-Laughlin chiral spin liquid (CSL) is one of the earliest and most important examples of a topologically-ordered quantum state [1, 2, 3] and has been extensively studied in the context of spin models [4, 5, 6, 7]. Recently, considerable effort has been devoted to identifying electronic settings in which the CSL may emerge [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. These studies set the stage for the investigation of charge fluctuation-driven transitions out of the CSL phase into other phases such as metals [8], superconductors [20], exotic charge-density waves [21], and quantum Hall states [22, 19].
@@
Figure 1: (a) Upper panel: The triangular Hofstadter-Hubbard model, with magnetic translations $T_{x,y}$, flux per triangle $\Phi_{\triangle}$, and Hubbard interaction $U$ indicated. Lower left panel: “butterfly” representation of non-interacting spectrum in the plane of $\Phi_{\triangle}$ and energy $E$. The bands at $\Phi_{\triangle}=\pi/2$ enjoy a doubled “magnetic” unit cell and particle-hole symmetry $\mathcal{P}$ relating opposite $C=\pm 1$ bands [lower right panel]. (b) Finite segment of the circumference-$L_{y}$ cylinder threaded by $\Phi_{\text{ext}}$ external flux, where $T_{y}$ translates around the cylinder and $x\in\mathbb{Z}$ indexes rings. (c) 2+1D phase diagram at a density of $n=1$ electron per site as a function of $U/t$ showing the integer quantum Hall (IQH) and chiral spin liquid (CSL) phases, with their transition marked by a star.
@@
The triangular Hofstadter-Hubbard model is thought to host both conventional and CSL electronic phases in the regime of strong orbital magnetic flux [23, 24, 6, 25, 26, 27, 28]. It is parameterized by an on-site interaction $U$ and hoppings of magnitude $t$ and phases consistent with a magnetic flux of $\Phi_{\triangle}$ per triangular plaquette (see Fig. 1a). When $\Phi_{\triangle}=\pi/2$, this system realizes a spin-singlet integer quantum Hall (IQH) insulator at small $U$ and density of one electron per site, with $\sigma_{xy}=2e^{2}/h$. At very large $U$, the state is a non-topological Mott insulator with a large charge gap, $120^{\circ}$ antiferromagnetic (AF) order, and gapless spin excitations described by a nearest-neighbor Heisenberg exchange of magnitude $\sim t^{2}/U$ [25, 26, 27, 28].
@@
A qualitative argument for the existence of an intermediate CSL phase is that the leading correction to the large-$U$ effective Heisenberg spin Hamiltonian is a chiral term $S\cdot S\times S$ with coefficient $J_{\triangle}\sim t^{3}\sin\Phi_{\triangle}/U^{2}$, which explicitly breaks both time-reversal and parity symmetry and favors a CSL state [24, 6, 29, 25, 26, 27, 28]. Very recently, Kuhlenkamp et al. presented cylinder iDMRG evidence of this intermediate CSL phase starting from $U\gtrsim 8.5t$ at flux $\Phi_{\triangle}=\pi/3$ and $U\gtrsim 10.5t$ at $\Phi_{\triangle}=\pi/2$ [19].
@@
Figure 2: CSL symmetry breaking at odd circumference. (a) The particle-hole operation $\mathcal{P}$ transforms the internal gauge flux through an odd-$L_{y}$ cylinder ring from $\phi$ to $\pi-\phi$. (b) Upper panel: In an odd-$L_{y}$ CSL ground state, the alternating pattern $(\pi/2,-\pi/2,\dots)$ of these internal fluxes at strong coupling spontaneously breaks $\mathcal{P}T_{x}$ symmetry, whose action is then equivalent to threading $\pi$ internal flux. Lower panel: the low-lying entanglement eigenvalues on the $L_{y}=5$ cylinder at spatial cuts between adjacent pairs of rings, which not only exhibit the approximate degeneracy expected for the CSL at each fixed $S_{z}$ [circled numerals], but also signal spontaneous breaking of $\mathcal{P}T_{x}$. (c) The internal flux configuration [upper panel] and low-lying entanglement eigenvalues [lower panel] of the IQH at $L_{y}=5$. The entanglement spectra at the two spatial cuts are equivalent.
The nature of the IQH-CSL transition, however, is not well understood. Parton mean field theories have been proposed for square lattice and honeycomb systems [22, 30]. Beyond this, there has been very little direct investigation into the nature of the putative critical point from either theory or numerics. In particular, the iDMRG investigation of Ref. [19] unearthed the IQH-CSL transition only as a crossover indicated by a weak maximum in a correlation length.
In this work, we use a compactification to infinite-length, finite-circumference cylinders (see Fig. 1b) in combination with a symmetry analysis and large-scale iDMRG simulations [31, 32, 33] to characterize the IQH and CSL states, as well as the transition between them. We show that a combination of particle-hole interchange and translation is spontaneously broken on odd-circumference cylinders in the CSL phase and present evidence that the IQH-CSL transition is in fact continuous in these geometries, with no intervening phase. We argue that the critical fluctuations are spin-singlet, with both spin and electron correlation functions decaying exponentially with distance. Considering also even-circumference cylinders, we argue that in 2+1D the IQH and CSL phases are separated by a continuous quantum phase transition that includes critical current fluctuations.
Figure 3: Evidence for the IQH-CSL continuous phase transition. (a) Upper panel: Sketch of two cylinder rings, illustrating the $\mathbb{Z}_{2}$ order parameter $\mathcal{O}_{0}$ defined as the total circumferential current operator [orange arrows]. Lower panel: $\langle\mathcal{O}_{0}\rangle$ versus $U/t$ for the odd-circumference cylinders, with $\langle\mathcal{O}_{0}\rangle\neq 0$ implying $\mathcal{P}T_{x}$ is spontaneously broken. (b-d) MPS transfer matrix correlation length in the $(Q,S_{z})=0$ charge sector for various MPS bond dimensions [increasing light to dark] and extrapolated to infinite bond dimension for $L_{y}=4$ [gray points]; the insets specify the external magnetic flux through the cylinder rings.
@@
Model, particle-hole symmetry and compactification—We study the Hofstadter-Hubbard model on the triangular lattice, defined pictorially in Fig. 1(a). In this model, electrons hop between nearest-neighboring sites and experience an Aharonov-Bohm phase of $\exp(i\Phi_{\triangle})$ upon orbiting a triangular plaquette (see App. A for details). We focus on the consequences of the orbital motion, setting the Zeeman coupling to zero [18, 19]. We fix $\Phi_{\triangle}=\pi/2$, which endows the Hamiltonian with a unitary particle-hole (PH) symmetry $\mathcal{P}$ chosen to act as
@@
(the spin rotation $i\sigma_{y}$ is included in $\mathcal{P}$ so it leaves the spin operator $\bm{S}=\frac{1}{2}c^{\dagger}\bm{\sigma}c$ invariant; see App. C). Denoting translation along the two non-parallel directions shown in Fig. 1 as $T_{x,y}$, which we represent in a gauge where
$T_{x}c^{\dagger}_{x,y}T_{x}^{-1}=(-1)^{y}c^{\dagger}_{x+1,y},\quad T_{y}c^{\dagger}_{x,y}T_{y}^{-1}=c^{\dagger}_{x,y+1},$ (2)
@@
we obtain the required magnetic translation algebra $T_{x}T_{y}=(-1)^{N_{F}}T_{y}T_{x}$, where $N_{F}$ is the fermion number. The commuting translations $(T_{x})^{2},T_{y}$ define a larger “magnetic” unit cell, yielding two sub-bands with Chern numbers $C=\pm 1$ related by $\mathcal{P}$ (Fig. 1a). We fix the filling to $n=1$ electron per lattice site, corresponding to full filling of the lowest spin-degenerate magnetic sub-band.
Dimensional reduction can break the $T_{x}$ symmetry. Consider the closed “tin can” shaped surface demarcated by the pair of adjacent lattice rings at $x$ and $x+1$. Applying Gauss’s law on this surface [34, 35], with $\Phi_{x}$ denoting the external flux through the cylinder ring at $x$, we have $\Phi_{x+1}+2L_{y}\Phi_{\Delta}\equiv\Phi_{x}$ modulo $2\pi$. The ring fluxes are therefore uniform when $L_{y}$ is even, but staggered when $L_{y}$ is odd. In the latter case, $T_{x}$ is explicitly broken.
Consider the odd circumference case and fix the ring fluxes to alternate between 0 and $\pi$. Though $T_{x}$ is broken, the combined translation-like operation $\mathcal{P}T_{x}$ (“glide-PH”) commutes with the Hamiltonian. The IQH state, being adiabatically connected to the unique ground state of the non-interacting Hamiltonian, respects $\mathcal{P}T_{x}$. However, the odd-$L_{y}$ CSL phase breaks $\mathcal{P}T_{x}$ spontaneously. To see this, note that in the CSL phase the charge fluctuations are gapped; when $U$ is large, the low energy degrees of freedom are described by an effective spin model [6]. Since $\mathcal{P}\bm{S}\mathcal{P}^{-1}=\bm{S}$, this spin model has a pure translation symmetry $t_{x}$. As each ring contains an odd number of spin-$1/2$ moments, the Lieb-Schulz-Mattis theorem requires that any gapped ground state be at least twofold degenerate [36], e.g., by breaking $t_{x}$ spontaneously. The same conclusion can be drawn from the anyon content of the CSL: the presence of one semion per unit cell [37, 38, 39, 34] causes $t_{x}$ to permute the two minimally-entangled ground states [40, 41, 42].
We can also interpret the CSL symmetry breaking in terms of the parton decomposition of the electron into a bosonic chargon $b$ carrying the charge and a fermionic spinon $f$ carrying the spin, $c_{\sigma}=bf_{\sigma}$ [43, 44, 45]. This requires introducing an internal gauge field $a$. In both the CSL and IQH phases, the saddle point configuration of the internal gauge field yields an internal flux through the triangular surface plaquettes that matches the external flux, $\phi_{\triangle}=\pi/2$ [5, 6, 20]. Moreover, assuming $L_{y}$ is odd, the internal flux through adjacent cylinder rings may be written as $(\phi,\phi+\pi)$, again differing by $\pi$ due to Gauss’s law. Since the CSL ground states are adiabatically connected to the Mott limit where $\mathcal{P}$ must act as the identity operator, at strong coupling their internal ring fluxes will energetically lock to $(\pi/2,-\pi/2)$ or $(-\pi/2,\pi/2)$, which are both invariant under the action of $\mathcal{P}$, namely $\phi\to\pi-\phi$ (Fig. 2a). These two configurations are related by $\mathcal{P}T_{x}\simeq t_{x}$, indicating spontaneous symmetry breaking (SSB) in these odd-$L_{y}$ geometries (Fig. 2b). In the IQH phase where the chargon is condensed, the Higgs mechanism instead pins together the internal and external gauge fields [46, 47, 48]. The resulting internal ring flux pattern $(0,\pi)$ respects $\mathcal{P}T_{x}$, consistent with a unique, gapped ground state (Fig. 2c, top).
Numerical observation of phase transition—We have used the iDMRG algorithm to determine the ground state on cylinders of circumference $L_{y}=3,4,5,6$. For the numerically studied range $U\leq 14$, we find no sign of the strong-coupling $120^{\circ}$ AF, so we focus on the IQH and CSL phases. We study the entanglement spectrum (ES) [49, 50] at a spatial bipartition of the system into left and right semi-infinite cylinders [8, 19]. The low-lying part of the ES corresponds to the conformal field theory (CFT) of the physical edge and exhibits degeneracies characteristic of the edge CFT [51, 50, 52, 53, 54, 55, 56, 57].
Consistent with the above expectations, the CSL phase spontaneously breaks $\mathcal{P}T_{x}$ symmetry on odd-$L_{y}$ cylinders: the low-lying entanglement eigenstates carry half-integer (integer) spin on bipartitions made to the right of even (odd) cylinder rings $x$ (see App. D for details). At the even cuts, the number of nearly-degenerate states with fixed $S_{z}=\hbar/2$ are $(1,1,2,3,5,\dots)$ at consecutive momenta; the counting of $S_{z}=0$ states on the odd cuts is similarly $(1,1,2,\dots)$ (see Fig. 2b, bottom). This pattern is characteristic of the free boson CFT [58]. In contrast, the low-lying ES of the IQH features states with both integer and half-integer spin (App. D). Consistent with unbroken $\mathcal{P}T_{x}$ symmetry, the spectrum is equivalent on every bipartition (Fig. 2c, bottom). The approximate degeneracies at fixed $S_{z}$ are $(1,2,5,10,20\dots)$, corresponding to two copies of the free boson CFT [59, 60, 61, 62].
We also construct a $\mathbb{Z}_{2}$ order parameter for $\mathcal{P}T_{x}$, $\mathcal{O}=\sum_{x\in\mathbb{Z}}J^{\perp}_{x}$ where $J^{\perp}_{x}$ is the average current on the circumferential bonds at $x$. In the chosen gauge at $\Phi_{\mathrm{ext}}=0$:
$J^{\perp}_{x}=(-1)^{x}L^{-1}_{y}\sum_{y}i(c_{x,y}^{\dagger}c_{x,y+1}-\mathrm{h.c.}).$ (3)
Since $(\mathcal{P}T_{x})J^{\perp}_{x}(\mathcal{P}T_{x})^{\dagger}=-J^{\perp}_{x+1}$, then $\mathcal{O}$ is charged under $\mathcal{P}T_{x}$. Since the ground states are invariant under $(\mathcal{P}T_{x})^{2}$ so that $\langle J^{\perp}_{x}\rangle=\langle J^{\perp}_{x+2}\rangle$, it suffices to examine the condensation of $\mathcal{O}_{0}=J^{\perp}_{0}+J^{\perp}_{1}$. In the CSL phase, $\langle\mathcal{O}_{0}\rangle\neq 0$ (see Fig. 3a), signaling SSB at $L_{y}=3,5$. Near the transition at odd $L_{y}$, the transfer matrix spectrum [63] contains a diverging correlation length in the $(Q,S_{z},k_{y})=0$ charge sector (see Fig. 3(b) for $L_{y}=5$ and App. G for $L_{y}=3$) consistent with the charge of the order parameter $\mathcal{O}$. Threading $\Phi_{\text{ext}}=\pi/2$ external flux through the odd-$L_{y}$ cylinder reduces the transition to a crossover since glide-PH symmetry is explicitly broken when $\Phi_{\text{ext}}\not\in\pi\mathbb{Z}$.
We further measure $\mathcal{O}_{\delta\mathrm{Heis}}=\sum_{m\in\mathbb{Z}}\mathcal{O}^{\delta\mathrm{Heis}}_{2m}$, defined by
@@
The order parameter $\mathcal{O}_{\delta\mathrm{Heis}}$ is likewise charged under $\mathcal{P}T_{x}$ and condenses in the CSL phase for odd $L_{y}$ and $\Phi_{\text{ext}}=0$ (see App. E). The two order parameters share the same symmetries and would be linearly coupled in a Ginzburg-Landau theory. We prefer $\mathcal{O}$ defined in Eq. (3) because it is larger in magnitude and simpler to express.